OpenStudy (mathlegend):

Example of a number written in fraction form that is not a rational number.

OpenStudy (mathlegend):

I thought any fraction was a rational number

OpenStudy (ghazi):

22/7 is in fractional form but not a rational number

OpenStudy (mathlegend):

oh pi... wow... I should have known. Thanks!

OpenStudy (ghazi):

:) YW

hero (hero):

\[\frac{e}{\pi}\]

hero (hero):

What about this question @MathLegend ?

OpenStudy (mathlegend):

Oh yeah.. but I do not know what e/π means

hero (hero):

e is the natural number and π is pi, neither of which are integers. By definition, a rational number is of the form \[\frac{a}{b}\] where both a and b are integers.

hero (hero):

By this definition, \(\frac{22}{7}\) is rational. It is an approximation for pi, however the numbers 22 and 7 are still integers.

hero (hero):

The only way you can make a fraction that is not rational is by using numbers that are not integers

hero (hero):

I just wanted to clear up any confusion in that regard